# Lessons from the high-resolution spectroscopy of AW UMa and $\epsilon\,$CrA: Is the Lucy model valid?

### v4

@@ 2 The two binaries as seen in radial velocities
    In both stars, the upper two-thirds of the primary component profile is narrower than predicted by the Lucy contact-binary model but agrees very well with a standard profile for a single star of a smaller radius rotating at the orbit-synchronous rotation rate (sizes 88% and 85% of the critical Roche-lobe side radius for AW UMa and $\epsilon\,$CrA, respectively). This property was recognized in AW UMa already by Rensing et al. (1985) using photographic spectra. The base of the primary profile in both binaries, referred to as the “pedestal” in P1, is roughly as wide, as predicted by the Lucy model. The narrow primary profile may be due to the anticyclonic circulation around the rotation pole of the mass-losing primary, as predicted by the hydrodynamic model of Oka et al. (2002), while the wide pedestal is explained by the binary circulation model of Stȩpień (2009). These observational properties contribute to our discussion of an appropriate model for W UMa binaries in Section 7.
@@
    In both stars, the primary $V\sin i$ widths show a tidal (ellipsoidal) component of $<\!5$ km s-1, which is very small in comparison to the mean values of $\simeq 181$ km s-1 for AW UMa (P1, Fig. 3) and $\simeq 147$ km s-1 for $\epsilon\,$CrA (P2, Fig. 3). These tidal variations are partly masked by greater noise in AW UMa but are better defined for $\epsilon\,$CrA.
@@
    The primary component of AW UMa showed a network of “ripples” or non-radial pulsations partly extending into the profile pedestal. The ripples, which are visible after subtraction of the stable mean profile (Figures 8, 9, 10 in P1), look very similar to the non-radial pulsations detected (subsequently to the work reported in P1) in the rapidly rotating, bright star Altair ($\alpha$ Aql, A7V, $V\sin i\simeq 240$ km s-1; Rieutord et al. (2023)).
    The primary of $\epsilon\,$CrA appears to be entirely free of such non-radial pulsations; they would be easily detectable because of the higher $S/N$ of its spectra.
@@
    The secondary-component RV profiles of both binaries appeared much weaker and flatter than expected by the Lucy model. This could indicate a genuine faintness of the secondaries or a different (later) spectrum, not matching the template used in the BF determination.
@@
    The secondary-component profiles in both binaries were distorted by localized perturbations which systematically shifted with the orbital phase. The disturbances, referred to as “wisps” in P1 and P2 due to their visibility in time-sequences of the RV profiles, complicated our interpretation of the secondary component shape in AW UMa; see their confusing appearance in Figure 13 in P1.
    $\epsilon\,$CrA (P2) showed only one “wisp”. Its phase evolution could be interpreted as due to a localized region of hotter (i.e.  primary-component) gas overlying the secondary component profile. Its location was estimated at the sub-observer orbital phase $\phi=0.65$ and may correspond to a collision front of the gas striking the primary after circulating the secondary component. The disturbance was very stable and remained visible during the whole observing run of the 28 orbital revolutions of $\epsilon\,$CrA. It is marked by an arrow in the $\phi=0.75$ profile in Figure 1; its orbital-phase evolution is discussed in Section 3. In this paper, we introduce the term Enhanced Spectral-line Perturbations (ESP) to stress that such perturbations in Broadening Function profiles correspond to stronger absorption spectra due to more gas atoms having the same atomic excitation properties as assumed for the early-F type template.
@@ 3 The secondary component of $\epsilon\,$CrA
The profiles were significantly fainter and flatter than those expected by the Lucy model (Figure 1). However, the extremes of the observed profiles were consistently well-defined to assume that their mean values could correspond to the motion of the secondary mass center. These values varied in anti-phase to the orbital motion of the primary component, allowing the derivation of the semi-amplitude $K_{2}$. This finding stands in starck contrast to the complex appearance of the secondary-component profiles in AW UMa (P1, Figure 13)). However, the differences in data quality and quantity between the studies of $\epsilon\,$CrA and AW UMa could influence this comparison, potentially favoring $\epsilon\,$CrA due to its better data quality and phase coverage.
@@ 4 The secondary component of AW UMa
Figure 6: The heliocentric orbital velocities of both AW UMa components. The figure corresponds to Figure 4 in P1, with added velocities of the secondary component, derived as the mean values of the secondary profile edges $E_{1}$ and $E_{2}$. The scale for the secondary (the right vertical axis) is compressed 10 times relative to that for the primary and is shifted to the same $V_{0}$. The three observational nights of AW UMa are coded by color: 1 – red, 2 – green, 3 – blue. The continuous lines give the sine fits to the data in the phase ranges as discussed in the text, with the solid-line portions marking the phase intervals used in the orbital solutions (see the text). The lower panel shows the half-width of the secondary component profile, $W_{2}=(E_{2}-E_{1})/2$, which for a standard rotational profile would correspond to the apparent equatorial velocity $V_{2}\sin i$.
@@ 5 The mass-ratio discrepancy
There is a discrepancy in the mass-ratio determinations for AW UMa, with spectroscopic results ($q_{\rm sp}$) consistently being larger than photometric results ($q_{\rm ph}$). The new spectroscopic determination, as listed in Table 3, $q_{\rm sp}=K_{1}/K_{2}=0.092\pm 0.007$, is close to the previous David Dunlap Observatory medium-resolution determination by Pribulla & Rucinski (2008), $q_{\rm sp}=0.101\pm 0.006$. The difference between the two determinations most likely reflects not only the use of different spectral resolutions but also the very different temporal coverage during the two observing programs; the DDO result utilized 10 partial nights distributed over a longer interval in contrast to the short, intense, hight-resolution run. It should be stressed that both determinations are larger than several photometric determinations by different authors (see references in P1, particularly Wilson (2008)), some claiming errors as small as 0.0005 tending to concentrate around $q_{\rm ph}=0.080\pm 0.005$. All photometric determinations utilized the Lucy model. As we discuss further in the paper, we are very sceptical about photometric determinations of $q_{\rm ph}$ utilizing the Roche/Lucy model for W UMa binaries, both due to the instability of the determination process (Section 6) and the physical viability of that model (Section 7).
@@
AW UMa was one of the first objects used to test modern spectroscopic techniques that utilized combined RV information from many spectral lines. This was mainly due to its brightness but also because of its unexpectedly small mass ratio indicated by the light-curve photometric solutions, $q_{\rm ph}\simeq 0.08$. Initially, the new spectroscopic approaches utilized the inherently non-linear Cross-Correlation Function (CCF) technique applied to also non-linear photographic data (McLean, 1981; Anderson et al., 1983; Rensing et al., 1985). These were later replaced by tests of a linear deconvolution (Broadening Function) technique applied to digital results (Rucinski, 1992). Because of low detector sensitivity, limited large telescopes time, and the unexpectedly weak signature of the secondary component in AW UMa, early analyses were limited to confirming the photometric mass ratio. The first spectroscopic determination specifically avoiding any prior assumptions and based on more extensive material by Pribulla & Rucinski (2008), $q_{\rm sp}=0.101\pm 0.006$, clearly demonstrated that the mass ratio must be larger than previously thought. Interestingly, this result coincided with a theoretical analysis of Paczyński et al. (2007), which pointed out difficulties in explaining AW UMa as the result of the close binary-star evolution for $q<0.10$.
@@
While the $q_{\rm sp}$ vs. $q_{\rm ph}$ discrepancy appears to be well established for AW UMa, a similar discrepancy is very weakly evident for $\epsilon\,$CrA, though still in the same direction, with $q_{\rm sp}=0.1300\pm 0.0010$ (P2) compared with $q_{\rm ph}=0.114\pm 0.003$ (Twigg, 1979) and $q_{\rm ph}=0.1244\pm 0.0014$ (Wilson & Raichur, 2011). It is not clear that $\epsilon\,$CrA exhibits total eclipses, which means that any photometric determination for that binary is less reliable than for a totally eclipsing binary such as AW UMa.
A comparison of $q_{\rm ph}$ with $q_{\rm sp}$ utilizing much better data than those available to Niarchos & Duerbeck (1991) is now possible. A large body of spectroscopic determinations resulted from a program of RV observations of short-period ($P<1$ day), bright binaries conducted in years 1993 – 2010 at the DDO and published in 23 papers in years 1993 – 2010. The program included 124 W UMa-type binaries. The spectroscopic determinations during the DDO program did not use any previous, literature data by design, ensuring no bias from published $q_{\rm ph}$; the binaries were selected for observations based simply on the sky accessibility, star brightness, and the somewhat erratic weather conditions at the DDO. A convenient list of the observed binaries is provided in Rucinski et al. (2013a, Table 1).
@@
The difference $(q_{\rm ph}-q_{\rm sp})$ is shown in Figure 8 directly and, similarly as in Niarchos & Duerbeck (1991), in the relative sense, divided by the value of $q_{\rm sp}$. The normalization may artificially amplify discrepancies at the low-end of the mass ratio. Regardless of the presentation, we observe a widespread occurrence of $q_{\rm ph}<q_{\rm sp}$, exactly as for AW UMa, which is the binary with the smallest $q_{\rm sp}$ in the figure. The size of the discrepancy is moderate, so either $q_{\rm sp}$ or $q_{\rm ph}$ could serve as an argument in the figure. However, we believe that the spectroscopic $q_{\rm sp}$ is currently the better-determined value because of its more astrophysically meaningful and straightforward derivation; we return to this subject in Section 7.
@@
The study of Niarchos & Duerbeck (1991) presented another strong indication that $q_{\rm ph}$ may be affected by phenomena occurring within the binary system, potentially affecting the applicability of the photometric Lucy model. The authors were confronted with excellent light curves of AE Phe that displayed significant season-to-season changes. One light curve showed a very well-defined total occultation eclipse, while the other light curves showed partial eclipses, usually associated with extensive out-of-eclipse light curve variability. This transition from the total to the partial eclipse already suggests a potential problem with the common-equipotential model for the binary. The authors modeled the four seasonal curves using the Lucy model and found a different $q_{\rm ph}$ for each season, with the range from $0.3836\pm 0.0014$ to $0.3964\pm 0.0010$. Although the formal uncertainties were small, as is almost always the case with the Lucy model solutions, a difference of more than 0.01 between individual determinations of $q_{\rm ph}$ is highly concerning. Strictly speaking, it would imply a large and rapid mass exchange between the components reaching $0.01\,M_{\odot}$ – but without an associated significant orbital-period jump. The Stȩpień model (Section 7.2) can explain the switches from the rounded to the flat-bottomed occultation eclipses in AE Phe as caused by variations in the overall loading of the flow (also called “equatorial bulge”) around the secondary component. Unfortunately, no reliable $q_{\rm sp}$ determination is available for that star at this moment.
@@ 6 The photometric and spectroscopic mass-ratio determinations
The tendency for $q_{\rm ph}<q_{\rm sp}$ is an important indication that one of the methods used for mass-ratio determinations – either the photometric Lucy model or the spectroscopic direct approach – may produce systematically incorrect results. We note that the results of this study, along with those from our two previous papers, definitely confirm the existence of a velocity field in the rotating system of coordinates. This fact eliminates the concept of equipotentials, which simply cannot be defined. In Section 7, we discuss the Stȩpień (2009) model for W UMa binaries, which explicitly includes velocities. This model is more complex than the Lucy model and necessitates abandoning the photometric determination of $q_{\rm ph}$. Before that, we will first consider the currently used photometric method of $q_{\rm ph}$ derivation, which does utilize equipotentials.
The one-to-one dependence of the Roche geometry on the mass ratio is a fundamental feature of the Lucy model and the foundation for its widespread success in light-curve synthesis codes. The dependence of the radius ratio, $r_{2}/r_{1}$, on the mass ratio, $q=M_{2}/M_{1}$, is a strict relation for an equipotential surface selected to represent the common envelope of the binary. We suggest that this assumption (1) likely leads to larger uncertainties in $q_{\rm ph}$ than claimed, and (2) may indicate that the $q_{\rm ph}$ vs. $q_{\rm sp}$ discrepancy highlights a failure of the Lucy model.
@@
The determination of photometric parameters for a W UMa-type binary is a complex process. Initially, a $3D\rightarrow 1D$ mapping of the stellar brightness (but not velocities needed for probing the dynamics) is performed by the revolving binary itself, which is registered as a light curve. This light curve is then subject to a multi-parameter solution utilizing the Roche-model geometry, with additional Lucy-model prescriptions for handling stellar-atmosphere properties. The solution is a highly non-linear, iterative process that includes unknowns such as the degree of contact (or the actual value of the common equipotential), and the orbital inclination, which are intermixed with assumed atmospheric properties, such as the local flux level estimated by surface brightness (effective temperature) and the limb and gravity darkening laws, typically derived from spherical stars. Due to inherent nonlinearities, this parameter determination process is highly susceptible to biased estimates and may result in deceptively small uncertainties when estimated as if for a linear, correlation-free problem based on the shape of the adopted deviations minimum. Unfortunately, this can lead to popularly quoted but suspiciously small formal uncertainties.
@@
The complexity of photometric determination of $q_{\rm ph}$ can contribute to the observed $q_{\rm ph}$ vs. $q_{\rm sp}$ discrepancy. To illustrate a prevailing nonlinearity, let’s specifically consider the size vs.  mass ratio dependence, which is essential for the $q_{\rm ph}$ determination. Eclipse effects carry relatively more information than the strongly model-dependent stellar-distortion effects. For spherical stars, it is known that the radius ratio, $k=r_{2}/r_{1}$ (Russell, 1912), is the key parameter controlling the depth and shape of eclipses. Although the binary components in our case are not spherical stars, an equivalent of $k$ exists, albeit hidden within complexities of light-curve-synthesis codes. The individual dimensions of Roche common equipotential surfaces, expressed, for example, as side equatorial radii $r_{1}$ and $r_{2}$, depend solely on the mass ratio, $r=r(q)$. This relation is known to be non-linear. A simplified version of it has been extensively used by practitioners of close-binary evolutionary calculations to estimate the size of the Roche-lobe (inner critical equipotential) filling component. In the orbital separation units, $r=0.38+0.2\log(q)$; this equation works for both components (with inverted $q$) and has an asymptotic low mass-ratio extension (Paczyński, 1971, Eq.4). By being approximately logarithmic, the dependence $r=r(q)$ is a slowly-varying function. The weak dependence of the Roche-lobe radii on $q$ is convenient for stellar binary models. However, $q_{\rm ph}$ determination utilizes an inverse relation, which – consequently – is a rapidly varying function. For the approximation as above, it is an exponential function, $q=10^{5r-1.9}$. Therefore, small errors in the relative radius determination produce large deviations in the estimated mass ratio. An error of 1% in $r$ produces a 4% error in an $q_{\rm ph}$ estimate.
@@
The non-linearity in the $q_{\rm ph}$ determination, as described above, has not yet been recognized. The use of different photometric codes and specific assumptions could potentially diminish its influence; thus, multiple independent solutions utilizing different codes might somehow average out the non-linearities. A more concerning issue is the tendency for $q_{\rm ph}<q_{\rm sp}$ among individual binaries, with some binaries exhibiting this behavior while others do not. While we strongly believe that an incorrect photometric model is the reason, some photometric solutions may provide systematically smaller values of $q_{\rm ph}$ due to the presence of an unrecognized “third light” in the light curve. Companions to W UMa binaries are very common, as demonstrated by Tokovinin et al. (2006) and Rucinski et al. (2007). A light curve containing unrecognized additional light has a reduced amplitude and may result in a $q_{\rm ph}$ value that is too small. That was the case for the photometric solution of one of our two targets, $\epsilon\,$CrA, analyzed by Shobbrook & Zola (2006); they commented on difficulties in obtaining the best fit starting from the assumed (and correct) spectroscopic value of $q_{\rm sp}=0.129$ (Goecking & Duerbeck, 1993), being forced to reduce $q_{\rm ph}$ by 0.02. $\epsilon\,$CrA is not listed in Table 4 precisely because it does not show total eclipses and therefore cannot provide a secure radius-ratio determination.
@@
The spectroscopic approach offers an entirely different and independent route of the mass ratio determination involving a $3D\rightarrow 2D$ mapping, from the spatial velocities into a series of RV profiles arranged in time. This approach utilizes more accessible information in two dimensions of the time and radial velocity; the RV data provide the orbital semi-amplitudes $K_{1}$ and $K_{2}$ leading to $q_{\rm sp}=K_{1}/K_{2}$. For very tight orbits, as observed for the W UMa-type binaries, with strongly dissipative, gaseous bodies involved, the only permitted orbits are circular, so that the derivation of the semi-amplitudes is particularly simple and direct. Thus – in principle – the spectroscopic $q_{\rm sp}$ is obtained with a lesser danger of biased results. Here, the main uncertainties relate to the proper localization of the RV centroids and their coincidence with the projected mass centers. Our results for AW UMa and $\epsilon\,$CrA in P1 and P2 show that the inter-binary velocity fields are present with velocities reaching perhaps as much as 30 – 50 km s-1. These fields may affect the observed velocity locations of the component mass centers. While the rotational-profile centroids seem to provide very well defined orbital velocities of the primary components of both of our targets, the confined velocity extent of the secondaries – similar to components of detached binaries – is an unexpected discovery and must be confirmed. It remains to be confirmed that centroids of such very faint features correctly trace the motion of the secondary components.

### v5

@@ 2 The two binaries as seen in radial velocities
    In both binaries, the upper two-thirds of the primary component profile is narrower than predicted by the Lucy contact-binary model but agrees very well with a standard profile for a single star of a smaller radius rotating at the orbit-synchronous rotation rate (sizes 88% and 85% of the critical Roche-lobe side radius for AW UMa and $\epsilon\,$CrA, respectively). This property was recognized in AW UMa already by Rensing et al. (1985) using photographic spectra. The base of the primary profile in both binaries, referred to as the “pedestal” in P1, is roughly as wide, as predicted by the Lucy model. The narrow primary profile may be due to the anticyclonic circulation around the rotation pole of the mass-losing primary, as predicted by the hydrodynamic model of Oka et al. (2002), while the wide pedestal is explained by the binary circulation model of Stȩpień (2009). These observational properties contribute to our discussion of an appropriate model for W UMa binaries in Section 7.
@@
    In both binaries, the primary-component profile widths show a small tidal distortion (line-of-centers elongation) of the more massive star of $<\!5$ km s-1. This number should be compared with the mean values of $V\sin i\simeq 181$ km s-1 for AW UMa (P1, Fig. 3) and $V\sin i\simeq 147$ km s-1 for $\epsilon\,$CrA (P2, Fig. 3). These width variations are partly masked by greater noise in AW UMa but are better defined for $\epsilon\,$CrA.
@@
    The primary component of AW UMa showed a network of “ripples” or non-radial pulsations partly extending into the profile pedestal. The ripples, which are visible after subtraction of the stable mean profile (Figures 8, 9, 10 in P1), look very similar to the non-radial pulsations detected (subsequently to the work reported in P1) in the rapidly rotating, bright star Altair ($\alpha$ Aql, A7V, $V\sin i\simeq 240$ km s-1 (Rieutord et al., 2023, 2024)) which is a $\delta$ Sct-type star. So far, there has been no attempt to detect photometrically these small, rapid pulsations in AW UMa. The primary of $\epsilon\,$CrA appears to be entirely free of non-radial pulsations; they would be easily detectable because of the higher $S/N$ of its spectra.
@@
    The secondary-component RV profiles of both binaries show significant deviations from the predictions of the Lucy model. They appear much weaker and flatter, which may indicate either a genuine faintness of the secondaries or a different (later) spectrum that does not match the template used in the BF determination.
@@
    The secondary-component profiles in both binaries were distorted by narrow perturbations of increased intensity of absorption lines. The disturbances, referred to as “wisps” in P1 and P2 due to their visibility in time sequences of the RV profiles, systematically shifted with the orbital phase, complicating our interpretation of the secondary component profile in AW UMa. See Figure 13 in P1 for an illustration of their confusing appearance. The wisps do not appear to result from the Struve-Sahade effect (Gies et al., 1997; Linder et al., 2007), which is observed in early-type stars. The S-S effect describes the increased strength of spectral lines of the approaching star over half of the orbital period. In contrast, the wisps in AW UMa and $\epsilon\,$CrA are narrow, well-localized in radial velocities, and their phase migratation allows their approximate geometric localization within the binary structure (Sections 3 & 4).
@@
    $\epsilon\,$CrA (P2) showed one well defined, strong, narrow “wisp”. However, an identical, strong feature appeared also in AW UMa – albeit among typically three to four wisps present there at a time. The phase evolution of this prominent wisp can be interpreted as due to a localized region of the primary-component gas overlying the secondary component profile. Its location in both stars was estimated at the sub-observer orbital phase $\phi=0.65$ and may correspond to a collision front of the gas striking the primary after circulating the secondary component. The disturbance was very stable and remained visible in $\epsilon\,$CrA during the whole observing run of the 28 orbital revolutions of the binary. It is marked by an arrow in the $\phi=0.75$ profile in Figure 1; its orbital-phase evolution is discussed in Section 3. In this paper, we introduce the term Enhanced Spectral-line Perturbations (ESP) to stress that such perturbations in Broadening Function profiles correspond to stronger absorption spectra due to spacially-localized gas atoms having the same atomic excitation properties as those of the early-F type primary star.
  7. 7.
@@ 3 The secondary component of $\epsilon\,$CrA
The profiles were significantly fainter and flatter than those expected by the Lucy model (Figure 1). However, the extremes of the observed profiles were consistently well-defined to assume that their mean values could correspond to the motion of the secondary mass center. These values varied in anti-phase to the orbital motion of the primary component, allowing the derivation of the semi-amplitude $K_{2}$. This finding stands in stark contrast to the complex appearance of the secondary-component profiles in AW UMa (P1, Figure 13). However, the differences in data quality and quantity between the studies of $\epsilon\,$CrA and AW UMa could influence this comparison, potentially favoring $\epsilon\,$CrA due to its better data quality and phase coverage.
@@ 4 The secondary component of AW UMa
Figure 6: The heliocentric orbital velocities of both AW UMa components. The figure corresponds to Figure 4 in P1, with added velocities of the secondary component, derived as the mean values of the secondary profile edges $E_{1}$ and $E_{2}$. The scale for the secondary (the right vertical axis) is compressed 10 times relative to that for the primary and is shifted to the same $V_{0}$. The three observational nights of AW UMa are coded by color: 1 – red, 2 – green, 3 – blue. The continuous lines show the sine curve fits to the data; the dashed line parts correspond to the phase intervals which were not used in the fits because of mutual interference of the wide component profiles. The lower panel shows the half-width of the secondary component profile, $W_{2}=(E_{2}-E_{1})/2$, which for a standard rotational profile would correspond to the apparent equatorial velocity $V_{2}\sin i$.
@@ 5 The mass-ratio discrepancy
There is a discrepancy in the mass-ratio determinations for AW UMa, with spectroscopic results ($q_{\rm sp}$) consistently being larger than photometric results ($q_{\rm ph}$). The new spectroscopic determination, as listed in Table 3, $q_{\rm sp}=K_{1}/K_{2}=0.092\pm 0.007$, is close to the previous David Dunlap Observatory medium-resolution determination by Pribulla & Rucinski (2008), $q_{\rm sp}=0.101\pm 0.006$. The difference between the two determinations most likely reflects not only the use of different spectral resolutions but also the very different temporal coverage during the two observing programs; the DDO result utilized 10 partial nights distributed over a longer interval in contrast to the short, intense, hight-resolution run. Both spectral determinations are larger than several photometric determinations by different authors (see references in P1, particularly Wilson (2008)), some claiming errors as small as 0.0005 tending to concentrate around $q_{\rm ph}=0.080\pm 0.005$.
@@
AW UMa was one of the first objects used to test modern spectroscopic techniques that utilized combined RV information from many spectral lines. This was mainly due to its brightness but also because of its unexpectedly small mass ratio indicated by the light-curve photometric solutions, $q_{\rm ph}\simeq 0.08$. First attempts utilized the inherently non-linear Cross-Correlation Function (CCF) technique applied to also non-linear photographic data (McLean, 1981; Anderson et al., 1983; Rensing et al., 1985). These were later replaced by a linear deconvolution (Broadening Function) technique applied to digital results (Rucinski, 1992). Because of low detector sensitivity, limited large telescopes time, and the unexpectedly weak signature of the secondary component in AW UMa, early analyses were limited to confirming the photometric mass ratio. The first spectroscopic determination based on more extensive material by Pribulla & Rucinski (2008), $q_{\rm sp}=0.101\pm 0.006$, clearly demonstrated that the mass ratio must be larger than previously thought. This result coincided with a theoretical analysis of Paczyński et al. (2007), which pointed out difficulties in explaining AW UMa as the result of the close binary-star evolution for $q<0.10$.
@@
A similar discrepancy as in AW UMa is very weakly evident for $\epsilon\,$CrA, though still in the same sense, with $q_{\rm sp}=0.1300\pm 0.0010$ (P2) compared with $q_{\rm ph}=0.114\pm 0.003$ (Twigg, 1979) and $q_{\rm ph}=0.1244\pm 0.0014$ (Wilson & Raichur, 2011). $\epsilon\,$CrA may show partial eclipses so that its photometric determination of $q_{\rm ph}$ may be less reliable than for a totally eclipsing binary such as AW UMa.
A comparison of $q_{\rm ph}$ with $q_{\rm sp}$ utilizing much better data than those available to Niarchos & Duerbeck (1991) is now possible. A large body of spectroscopic determinations resulted from a program of RV observations of short-period ($P<1$ day), bright binaries conducted in years 1993 – 2010 at the DDO and published in 23 papers in years 1993 – 2010. The program included 124 W UMa-type binaries. The spectroscopic determinations during the DDO program did not use any previous, literature data by design, ensuring no bias from published $q_{\rm ph}$. The binaries were selected for observations based on the sky accessibility, star brightness, and the somewhat erratic weather conditions at the DDO. A list of the observed binaries is provided in Rucinski et al. (2013a, Table 1).
@@
The difference $(q_{\rm ph}-q_{\rm sp})$ is shown in Figure 8 directly and, similarly as in Niarchos & Duerbeck (1991), in the relative sense, divided by the value of $q_{\rm sp}$. The normalization is kept here to reflect our lack of knowledge on the source of the discrepancy. Regardless of the presentation, we observe a widespread occurrence of $q_{\rm ph}<q_{\rm sp}$, exactly as for AW UMa, which is the binary with the smallest $q_{\rm sp}$ in the figure. The size of the discrepancy is moderate, so either $q_{\rm sp}$ or $q_{\rm ph}$ could serve as an argument in the figure. The author believes that the spectroscopic $q_{\rm sp}$ results are currently better-defined because of their more straightforward derivation. We return to this subject in Section 7.
@@
The study of Niarchos & Duerbeck (1991) provided another strong indication that $q_{\rm ph}$ may be affected by phenomena not accounted in the photometric model. The light curve of AE Phe showed significant season-to-season changes, with one light curve showing a well-defined total occultation eclipse. The authors modeled the four seasonal curves using the Lucy model and found a range of $q_{\rm ph}$ results, from $0.3836\pm 0.0014$ to $0.3964\pm 0.0010$. While the formal uncertainties are undoubtedly under-estimated, a difference of more than 0.01 between individual determinations of $q_{\rm ph}$ seems to be real and is highly concerning. Such a discrepancy would suggest a large and rapid mass exchange between the components, amounting to $0.01\,M_{\odot}$, yet without an accompanying significant orbital-period variation.
The Stȩpień model (Section 7.2) offers a possible explanation for the transitions between rounded and flat-bottomed occultation eclipses in AE Phe. These variations could result from changes in the overall loading of the flow (also referred to as the “equatorial bulge”) around the secondary component. Unfortunately, no reliable $q_{\rm sp}$ determination is available for this star at present.
@@ 6 The photometric and spectroscopic mass-ratio determinations
The tendency for $q_{\rm ph}<q_{\rm sp}$ is an important indication that one of the methods used for mass-ratio determinations – either the photometric Lucy model or the spectroscopic direct approach – may produce systematically incorrect results. We note that the results of this study, along with those from our two previous papers, definitely confirm the existence of a velocity field in the rotating system of coordinates. Thus, strictly speaking, equipotentials cannot be defined. However, for low velocities, stellar shapes may still be close to predicted to the equipotential geometry. We return to this subject in Section 7.
@@
The determination of photometric parameters for a W UMa-type binary is a complex process. Initially, a $3D\rightarrow 1D$ mapping of the stellar brightness (but not velocities) is performed by the revolving binary itself, which is registered as a light curve. This light curve is then subject to a multi-parameter solution utilizing the Roche-model geometry, with additional Lucy-model prescriptions for handling stellar-atmosphere properties. The solution is a highly non-linear, iterative process that includes unknowns such as the degree of contact (or the actual value of the common equipotential), and the orbital inclination, which are intermixed with assumed atmospheric properties, such as the local flux level estimated by surface brightness (effective temperature) and the limb and gravity darkening laws, typically derived from spherical stars. Due to inherent nonlinearities, this parameter determination process is highly susceptible to biased estimates and may result in deceptively small uncertainties when estimated as if for a linear, correlation-free problem based on the shape of the adopted deviations minimum. Unfortunately, this can lead to popularly quoted but suspiciously small formal uncertainties.
@@
The complexity of photometric determination of $q_{\rm ph}$ can contribute to the observed $q_{\rm ph}$ vs. $q_{\rm sp}$ discrepancy. Eclipse effects carry relatively more information than the strongly model-dependent stellar-distortion effects. For spherical stars, it is known that the radius ratio, $k=r_{2}/r_{1}$ (Russell, 1912), is the key parameter controlling the depth and shape of eclipses. Although the binary components in our case are not spherical stars, an equivalent of $k$ exists, albeit hidden within complexities of light-curve-synthesis codes. The individual dimensions of Roche common equipotential surfaces, expressed, for example, as side equatorial radii $r_{1}$ and $r_{2}$, depend solely on the mass ratio, $r=r(q)$. This relation is known to be non-linear. A simplified version of it has been extensively used by practitioners of close-binary evolutionary calculations to estimate the size of the Roche-lobe (inner critical equipotential) filling component. In the orbital separation units, $r=0.38+0.2\log(q)$; this equation works for both components (with inverted $q$) and has an asymptotic low mass-ratio extension (Paczyński, 1971, Eq.4). By being approximately logarithmic, the dependence $r=r(q)$ is a slowly-varying function. The weak dependence of the Roche-lobe radii on $q$ is convenient for stellar binary models. However, $q_{\rm ph}$ determination utilizes an inverse relation, which – consequently – is a rapidly varying function. For the approximation as above, it is an exponential function, $q=10^{5r-1.9}$. Therefore, small errors in the relative radius determination produce large deviations in the estimated mass ratio. An error of 1% in $r$ produces a 4% error in an $q_{\rm ph}$ estimate.
@@
The non-linearity in the $q_{\rm ph}$ determination, as described above, has not yet been recognized. A more concerning issue is the tendency for $q_{\rm ph}<q_{\rm sp}$ among individual binaries (Section 5), with some binaries exhibiting this behavior while others do not. While we believe that an incorrect photometric model is the reason, some photometric solutions may provide systematically smaller values of $q_{\rm ph}$ due to the presence of an unrecognized “third light” in the light curve. Companions to W UMa binaries are very common, as demonstrated by Tokovinin et al. (2006) and Rucinski et al. (2007). A light curve containing unrecognized additional light has a reduced amplitude and may result in a $q_{\rm ph}$ value that is too small. That was the case for the photometric solution of one of our two targets, $\epsilon\,$CrA, analyzed by Shobbrook & Zola (2006); they commented on difficulties in obtaining the best fit starting from the assumed (and correct) spectroscopic value of $q_{\rm sp}=0.129$ (Goecking & Duerbeck, 1993), being forced to reduce $q_{\rm ph}$ by 0.02. $\epsilon\,$CrA is not listed in Table 4 precisely because it does not show total eclipses and therefore cannot provide a secure radius-ratio determination.
@@
The spectroscopic approach offers an entirely different and independent route of the mass ratio determination involving a $3D\rightarrow 2D$ mapping, from the spatial velocities into a series of RV profiles arranged in time. This approach utilizes more accessible information in two dimensions of the time and radial velocity; the RV data provide the orbital semi-amplitudes $K_{1}$ and $K_{2}$ leading to $q_{\rm sp}=K_{1}/K_{2}$. For very tight orbits, as observed for the W UMa-type binaries, with strongly dissipative, gaseous bodies involved, the only permitted orbits are circular, so that the derivation of the semi-amplitudes is particularly simple.
Although the direct approach using velocities suggests a lesser risk of biased results for the spectroscopic $q_{\rm sp}$, the values of the amplitudes $K_{i}$ may still be biased for other reasons. Paradoxically, the $K_{2}$ amplitude, which was very hard to determine in our binaries, seems to be more reliable, but only under strict conditions on the spectral feature used for the $K_{2}$ estimate. The feature: (1) must be free of any component profile overlap, (2) must move with the orbital phase in a sinusoidal manner, (3) and strictly in anti-phase to the motion of the primary component. Obvioulsy, any inter-binary velocity features must be entirely avoided.
In contrast to $K_{2}$, the primary component velocity amplitude $K_{1}$ appeared deceptively easy to determine in both binaries analyzed here. The primary profiles seemed symmetric and straightforward to approximate using the classical rotational profile. However, it is unclear whether the profile center accurately traces the motion of the mass center, due to the distortion of the primary component. In both binaries, the small $v\sin i$ width variation due to the symmetric term of the simplfied Roche potential expansion, $J=2$ (e.g.  Eq.(2) in Rucinski (1969)) was estimated at $\leq\!5$ km s-1 (item #2 in the list in Section 2). The first asymmetric term ($J=3$) is expected to be smaller by a factor of roughly $r/a$ compared to the dominant tidal elongation term, where $r$ is the primary mean radius and $a$ is the mass center sepraration. The expected shift, $<\!1.5-2.0$ km s-1, would lead to a systematic error $\delta(q_{\rm sp})\simeq 0.007$. Notably, due to the sense of the Roche-lobe asymmetry, this error would be opposite to that suggested by small value of $q_{\rm ph}$, i.e. the actual $K_{1}$ of the mass centre (and thus $q_{\rm sp}$) would be larger than estimated for the symmetric profile.